# On the generalization of language models from in-context learning and finetuning: a controlled study

### v2

@@ Abstract
Large language models exhibit exciting capabilities, yet can show surprisingly narrow generalization from finetuning. E.g. they can fail to generalize to simple reversals of relations they are trained on, or fail to make simple logical deductions based on trained information. These failures to generalize from fine-tuning can hinder practical application of these models. On the other hand, language models’ in-context learning shows different inductive biases, and can generalize better in some cases. Here, we explore these differences in generalization between in-context- and fine-tuning-based learning. To do so, we constructed several novel datasets to evaluate and improve models’ abilities to generalize from finetuning data. The datasets are designed to create clean tests of generalization, by isolating the knowledge in the dataset from that in pretraining. We expose pretrained large models to controlled subsets of the information in these datasets — either in context, or through fine-tuning — and evaluate their performance on test sets that require various types of generalization. We find overall that in data-matched settings, in-context learning can generalize more flexibly than fine-tuning (though we also find some qualifications of prior findings, such as cases when fine-tuning can generalize to reversals embedded in a larger structure of knowledge). We build on these findings to propose a method to enable improved generalization from fine-tuning: adding in-context inferences to finetuning data. We show that this method improves generalization across various splits of our datasets and other benchmarks. Our results have implications for understanding the inductive biases of different modes of learning in language models, and practically improving their performance.
@@ Introduction
Language models (LMs) pretrained on huge corpora of internet text become efficient in-context learners; they can generalize from a few examples of a task to answer new instances (Brown et al., 2020; Gemini Team Google, 2023). Pretrained LMs can also be fine-tuned for downstream tasks using relatively few examples—though achieving good generalization from fine-tuning often requires hundreds to thousands of examples (e.g. Kirstain et al., 2022; Vieira et al., 2024). Indeed, the generalization from fine-tuning on a particular example can be surprisingly limited; for example, LMs fine-tuned on a statement like “B’s mother is A” fail to generalize to answer “Who is A’s son?” (Berglund et al., 2024; cf. Allen-Zhu and Li, 2025). However, LMs can readily answer questions about these types of reverse relations in context (e.g. Lampinen et al., 2024b). Furthermore, transformers show different inductive biases when generalizing from weights vs from context (cf. Chan et al., 2022b; Shen et al., 2023; Russin et al., ). How do patterns of generalization differe between in-context learning and fine-tuning? What are the implications for how we should adapt models for new tasks or information? In this paper, we explore these questions.
To do so, we construct controlled synthetic datasets of factual knowledge. We design these datasets to have complex and self-consistent structures, while avoiding any overlap with knowledge that may be present in pretraining corpora. We create train and test splits of these datasets that involve different types of generalization: reversals or chaining multiple logical inferences into a syllogism. We then evaluate how well large pretrained language models generalize to these test sets through fine-tuning, or through in-context learning—via placing the entire training set (or large subsets thereof) in context. We also explore various methods of improving generalization, such as data augmentation. Overall, we find that across various datasets, in-context learning (ICL) generalizes better than finetuning. However, finetuning generalization can be improved, and in fact can even exceed ICL generalization, by spending more train time compute to augment the training dataset using in-context inferences.
@@
  • We study the distinct patterns of generalization that pretrained LMs exhibit from in-context learning and finetuning.
  • We find that in-context learning over the entire training dataset often generalizes better than finetuning does, when evaluated on systematic holdouts like reversals, syllogistic inferences, compositions, etc.
  • We propose to bridge this gap via dataset augmentation—prompting an LM to generate augmentations of the dataset in-context, and then adding these augmented data to the training set.
@@
We first test our methods with two simple datasets [^1] These datasets are adapted from the nonsense NLI and Syllogism datasets from Lampinen et al. 2024b. that contain independent examples of reversal and syllogistic reasoning.
Simple reversals: Each training example consists of a brief preamble (e.g. “Did you know that”) followed by a single statement comparing two entities, e.g. “femp are more dangerous than glon.” There are a hundred such comparisons provided (with the comparison words sampled from a set of 28, e.g. “brighter”, “heavier,”). Each comparison is repeated across 10 different training examples, each time paired with a randomly sampled preamble. The test set consists of forced choices between the correct reversal and a contradictory relation, e.g.: “glon are less/more dangerous than femp.”
Simple syllogisms: There are 69 training examples. Each example consists of a preamble (which specifies the entities under discussion), followed by two statements that form a syllogism. For example: “The relationships between glon, troff and yomp are: All glon are yomp. All troff are glon.” The test examples evaluate whether the model makes a correct inference from that syllogistic form. They provide a preamble specifying two entities from the training example, and then score all possible statements about the relationship between the two entities involving the quantifiers and relations in the dataset. Following Lampinen et al. (2024b), we omit the partial negative form (some X are not Y); thus, there are six such possible statements (the product of “all”, “some”, and “no” together with the two directions each of those relations could take). For example, the preamble for the test example corresponding to the syllogism above would be “The relationships between yomp and troff are:” and the correct answer would be “All troff are yomp.”
Reversal curse paper
We use the reversal dataset proposed by Berglund et al. (2024) that contains one-sentence descriptions of fictional celebrities. Dataset examples can have the celebrity name (e.g. ‘Daphne Barrington’) preceding the description (e.g. ‘the director of "A Journey Through Time."’) or vice-versa. During finetuning, the training dataset is presented in one order, e.g. where the name precedes the description, and then tested on the reversed order, where the name follows the description.
Following Berglund et al. (2024), we use two independent sets of celebrities – ‘A’ and ‘B’. During finetuning, we present the celebrities from the set ‘A’ with names preceding the description, and from set ‘B’ where names and descriptions appear in both orders but in separate examples. Overall, the train set consists of 3600 examples. The test set evaluates whether the model can infer the celebrity names from the set ‘A’ given only their descriptions. To add distractors in test examples, we include three randomly selected incorrect celebrity names in the list of options to score.
@@
Figure 1: A semantic structure with a hierarchy of properties and relations. (Reproduced with permission from Rogers and McClelland (2008).)
The benchmark is built around a relational semantic hierarchy that allows deductive inferences and abstractions. This hierarchy is based on real-world categories and relations, and draws inspiration from prior work in cognitive science that has studied learning of rich semantic hierarchies (see Fig. 1 for an intuitive example that composes part of the real-world structure). We similarly create a hierarchical knowledge structure involving 110 categories of animals and objects, and properties (1-6 per category, plus inherited ones) and relations. This underlying structure is derived from the real world.
@@
The use of nonsense terms potentially poses tokenization challenges; however, we ameliorate these by generating short 4-5 letter nonsense words using plausible combinations of phonemes for English (sampled via Gao et al., 2023). Moreover, we show in the next sections that in practice the models can readily make inferences about these nonsensical entities in context, so use of unfamiliar words is not the primary bottleneck on model performance.
@@
Test sets: Within the semantic structure, our test sets accommodate reversals of relations (gruds [dogs] are a type of abmes [mammals] => abmes include gruds; cf. Berglund et al., 2024), syllogism-like deductive inferences (e.g. gruds [dogs] are abmes [mammals]; abmes are rony [warm-blooded] => gruds are rony), and longer deductions. Specifically, we focus on the following splits (in rough difficulty order):
@@
  • Syllogisms over trained facts.
  • Category holdouts: only one fact about a category is present in training: what its parent category is. All possible inferences from that fact are tested. This overlaps with some aspects of the syllogism splits, except that the information about the target category is more strictly limited, thus limiting other cues that could aid generalization.
@@ Methods
We perform evaluation using multiple-choice likelihood scoring. We do not provide the answer choices in context.
@@
Our fine-tuning experiments mainly involve tuning Gemini 1.5 Flash (Gemini Team Google, 2024) on our datasets. We generally fine-tune with batch size 8 or 16 and learning rate $3\cdot 10^{-4}$ for 200-1000 steps (depending on dataset size and loss).
In-context evaluation
To perform full-dataset in-context evaluation, we concatenate the documents in a training dataset and provide them as context to the (instruction-tuned) model. We then prompt the model to use that context to answer the question. We randomly subsample the documents by a factor of 8x when evaluating models in-context on the largest datasets, as there is some redundancy in the datasets, and we find that models experience interference as the context length scales.
@@
Our key approach to dataset augmentation is to take advantage of the in-context generalization of models in order to improve the coverage of the fine-tuning dataset. We approach this via several methods, but all chiefly aim at the goal of spending training time compute for in-context inference to create more finetuning data, in order to improve generalization out-of-context (i.e., without requiring additional information in context) at test time.
@@
Some datasets, such as the fictional celebrities dataset by Berglund et al. (2024) and our semantic structure dataset, contain documents comprising multiple sentences that are logically or semantically linked. We find that splitting these documents at sentence-level, into multiple finetuning examples, leads to a much improved finetuning performance – even after accounting for the total dataset size and gradient steps. We explore two ways of splitting a document: 1) Independent Splitting: wherein all $n$ sentences of a document are split independently into $n$ separate training examples, 2) Cumulative Splitting: wherein an $n$-sentence document is split into $n$ examples, with the $i$th example contains all sentences up to the $i$th sentence. We analyze the effect of sentence splitting on the model generalization in appendix B.1. In the following sections, we assume independent sentence splitting unless stated otherwise.
@@
In Fig. 2, we first explore generalization in the context of the reversal curse phenomenon and dataset released by Berglund et al. (2024). We replicate the finding that finetuning on the forward directions does not produce generalization to the reversals. The authors of that work noted briefly that in-context the models could perform better at this task. We study this more systematically by presenting the entire dataset in context, and find that the model performs nearly at ceiling on the reversals—thus providing a strong demonstration of the benefits of in-context learning over finetuning. Finetuning with data augmented with in-context inferences yields similarly high test performance. Simple finetuning, on the other hand, has near zero accuracy as the finetuned model always prefers those (incorrect) celebrity name completions that were seen as targets during training, regardless of the context. Finally, a pretrained model performs near chance on the test set, indicating a lack of contamination.
Simple nonsense reversals
We then correspondingly test the models on the simple nonsense version of our reversal dataset (Fig 3, left). We find a weaker, but still noticeable, benefit of ICL over finetuning in this setting, and again stronger benefits of augmented finetuning. The difference in benefits compared to the experiments above may be due to differences in the plausibility of the relationships in question, e.g. the possibility that nonsense words interfere to some degree with the model’s reasoning over longer contexts (see Appx. B.2).
Simple syllogisms
We next test the simple syllogisms dataset (Fig 3, right). Again, the pretrained model performs at chance, indicating a lack of contamination. Finetuning on the dataset does yield some non-trivial above-chance generalization; perhaps because for certain flavors of syllogism the logical inferences are compatible with simpler linguistic patterns — e.g., tuning on sequences like “All X include Y.” “All Y include Z.” might make the model more likely to predict “All X include Z” from pure associations. However, following simpler associative patterns is not valid for most syllogistic forms, e.g. if the universal quantifier is replaced by the existential in the examples above. Perhaps because of this, in-context learning yields much stronger performance; using in-context learning to augment the training dataset further improves performance overall.
@@
We then test the semantic structure benchmark (Fig. 4), which integrates multiple kinds of generalization in a richer learning setting. In this setting, we evaluate performance on: a) rephrased training statements (that preserve the direction of the relation) denoted as ‘train’ in the figure, b) reversals, c) syllogistic inferences, and d) propositions about held out categories. Across these settings, we find an overall benefit of ICL over finetuning, though the magnitude of this benefit varies depending on the split. We find some improvements in generalization to even rephrased information from training, and more dramatic improvements on reversals and syllogisms. However, the category-holdout split remains difficult, and improvements from ICL are minimal. Furthermore, we continue to find that augmenting the finetuning data with in-context inferences improves performance from finetuning, in many cases outperforming even ICL. (N.B., in this setting we do not use sentence splitting for the finetuning baseline, as we find it impairs performance and we want to compare to the strongest baselines; results with sentence-splitting for the baseline, along with other ablations, can be found in Appx. B.4.)
Our results in this section also highlight an important nuance to prior results on the reversal curse (Berglund et al., 2024). When the information being tested forms part of a broader coherent structure of knowledge (such as our semantic structure), finetuning alone does exhibit some above-chance generalization to reversals. This generalization may be due to the fact that other information in the training set can support the reversed conclusion; e.g. if the reversal is from “birds include eagles” to “eagles are a type of bird,” but the training set also includes statements “eagles have wings,” that information may enable an alternative route to inferring the reversal (associatively if not logically). Nevertheless, in-context learning and augmented finetuning continue to perform substantially better than finetuning alone.
@@
Process knowledge
We also performed some preliminary experiments exploring learning and generalization of process-type (rather than semantic) knowledge, i.e. the ability to apply novel procedures to inputs. Specifically, we focused on executing a simple pseudo-mathematical procedure that we call “derivatoids”, which transforms math-like expressions in a manner similar to computing a derivative. We created a train-test compositional generalization split in which certain combinations of rules are seen in training (again, either in context or via finetuning), but other combinations are held out for testing. The overall pattern of results is generally consistent with those above. ICL outperforms finetuning, especially in the low-shot regime, and augmentations improve finetuning performance. However, process knowledge requires distinct evaluation methods, and the effects correspondingly appear to be driven by distinct factors compared to the semantic knowledge experiments above; thus, we present the experiments and results in full detail in Appendix B.5.
@@
In-context learning: A variety of works have explored in-context learning (Brown et al., 2020; Dong et al., 2022; Lampinen et al., 2024a). Several works have studied patterns in the learning and generalization of ICL, either empirically (Garg et al., 2022; Chan et al., 2022a), mechanistically (Olsson et al., 2022), or theoretically (Xie et al., 2023; Zhang et al., 2024; Elmoznino et al., 2024). Several recent works have found that even scaling to hundreds or thousands of in-context examples can improve LM performance on challenging datasets (Agarwal et al., 2024; Anil et al., 2024; Bertsch et al., 2024). Our work similarly relies on the ability of models to learn from many documents in context in order to generate the inferences for augmentation.
@@
Out-of-context learning: Several other papers (e.g. Berglund et al., 2023; Meinke and Evans, 2023) have considered how models can generalize “out-of-context”—that is, how they can use information that is not directly included in the prompt in a flexible way at test time. Our results may connect to these, e.g. in the patterns of partial generalization seen even with finetuning on the semantic structure dataset. However, we generally do not find reliable use of out-of-context information as in-context—i.e., in-context learning tends to outperform finetuning.
@@
Factual learning & generalization: A variety of works have studied how language models learn and generalize factual information, both in pretraining (e.g. Allen-Zhu and Li, 2024, 2025; Zucchet et al., 2025) and during finetuning (e.g. Berglund et al., 2024; Gekhman et al., 2024). Many of these works have focused on failures to generalize like the “Reversal Curse” (Berglund et al., 2024), while others looked into increases in hallucination as LLMs are pretrained (Zucchet et al., 2025) or finetuned on new knowledge (Gekhman et al., 2024). However, these works have generally not performed data-matched experiments comparing the generalization of in-context learning and finetuning.
@@
Data augmentation: A wide variety of works have explored how LLMs can be used to augment data for improved performance from small or narrow datasets, e.g. to improve zero-shot task performance (Maini et al., 2024) or to improve cross-lingual generalization (e.g. Whitehouse et al., 2023). Ding et al. (2024) reviews some of this broader literature. There have also been targeted attempts to fix specific issues like the reversal curse with hardcoded augmentations (Golovneva et al., 2024). A closely related work by Akyürek et al. (2024) proposes “deductive closure training” to improve coverage via prompting language models to generate deductive inferences from training documents. Padmanabhan et al. (2024) similarly propose generating continuations and then distilling them into the model. Several concurrent works Chen et al. (2024); Ruan et al. (2025) suggest that having LMs generate additional reasoning directions, and using these to augment their training data, can improve performance on reasoning tasks. Yang et al. (2024) use language models to extract entities in the training documents, and generate synthetic data reasoning about links between these entities; like our results, they find this method of “rearranging” knowledge in synthetic data helps to improve downstream performance. Our experiments show that in controlled settings, without the possibility of dataset contamination, similar approaches to augmenting small finetuning datasets can yield improved generalization—and relate the performance to that achieved through more basic finetuning and in-context learning.
@@
Synthetic data: An equally broad literature has explored other applications of synthetic data to improving LM performance; see Liu et al. (2024) for a recent survey. Earlier works considered hand-designed synthetic data targeted to improve generalization using domain knowledge of areas like linguistics or mathematics (e.g. Pratapa et al., 2018; Wu et al., 2021). More recent approaches have focused on generating the data directly from language models, either filtering with ground-truth scores (Zelikman et al., 2022) or simply through self-consistency (Wang et al., 2023; Huang et al., 2023). While a recent prominent article argued that training on synthetic data leads models to collapse (Shumailov et al., 2024), other works have noted that in settings where synthetic data is mixed with the original data performance continues to improve (Gerstgrasser et al., 2024). We correspondingly find that in our setting, rather than being harmful, synthesizing augmented data with the models consistently improves performance (even on rephrased information from the training split in the semantic structure dataset). These results contribute to the ongoing discussion of how incorporating synthetic data affects model performance.

### v3

@@ Abstract
Large language models exhibit exciting capabilities, yet can show surprisingly narrow generalization [^1] By generalization we mean simple logical deductions of the trained knowledge which is the fundamental aspect of reasoning. from finetuning. E.g. they can fail to generalize to simple reversals of relations they are trained on, or fail to make simple logical deductions based on trained information. These failures to generalize factual information from fine-tuning can significantly hinder the reasoning capabilities of these models. On the other hand, language models’ in-context learning shows different inductive biases and deductive reasoning capabilities. Here, we explore these differences in generalization and deductive reasoning between in-context- and fine-tuning-based learning. To do so, we constructed several novel datasets to evaluate and improve models’ abilities to make generalizations over factual information from novel data. These datasets are designed to create clean tests of generalization, by isolating the knowledge in the dataset from that in pretraining. We expose pretrained large models to controlled subsets of the information in these datasets — either in context, or through fine-tuning — and evaluate their performance on test sets that require various types of generalization. We find overall that in data-matched settings, in-context learning can generalize several types of inferences more flexibly than fine-tuning (though we also find some qualifications of prior findings, such as cases when fine-tuning can generalize to reversals embedded in a larger structure of knowledge). We build on these findings to propose a method to enable improved generalization from fine-tuning: adding in-context reasoning traces to finetuning data. We show that this method improves generalization and deductive reasoning across various splits of our datasets and other benchmarks. Our results have implications for understanding the different forms of reasoning afforded by different modes of learning in language models, and practically improving their performance.
@@ Introduction
Language models (LMs) pretrained on huge corpora of internet text become efficient in-context learners; they can generalize from a few examples of a task to answer new instances (brown2020language; team2023gemini). Pretrained LMs can also be fine-tuned for downstream tasks using relatively few examples—though achieving good generalization from fine-tuning often requires hundreds, or even thousands of examples (e.g. kirstain2022few; vieira2024much). Indeed, the generalization from fine-tuning on a particular example can be surprisingly limited; for example, LMs fine-tuned on a statement like “B’s mother is A” fail to generalize to answer “Who is A’s son?” (berglund2024reversal; cf. allen2025physics3_2). These failures of simple logical deduction of finetuned knowledge can hinder reasoning capabilities of LMs. However, LMs can readily reason through questions about these types of reverse relations in context (e.g. lampinen2024language). Furthermore, transformers show different inductive biases when generalizing from weights vs. from context (cf. chan2022transformers; shen2023pretrained; russin2024human). How do patterns of generalization differ between in-context reasoning and fine-tuning? What are the implications for how we should adapt models for new tasks or information? In this paper, we explore these questions.
To do so, we construct controlled synthetic datasets of factual knowledge. We design these datasets to have complex and self-consistent structures, while avoiding any overlap with knowledge that may be present in pretraining corpora. We create train and test splits of these datasets that involve different types of generalization: reversals or chaining multiple logical inferences into a syllogism. We then evaluate how well large pretrained language models generalize to these test sets through fine-tuning, or through in-context learning—via placing the entire training set (or large subsets thereof) in context. These results show that models can use long contexts effectively to generalize better than gradient-based learning. We also explore various methods of improving generalization, such as data augmentation. Overall, we find that across various datasets, in-context learning (ICL) generalizes better than finetuning. However, finetuning generalization can be improved, and in fact can even exceed ICL generalization, by spending more train time compute to augment the training dataset using in-context reasoning.
@@
  • We study the distinct patterns of factual generalization that pretrained LMs exhibit from in-context learning and finetuning.
  • We find that in-context learning over the entire training dataset often generalizes better than finetuning, when evaluated on systematic holdouts like reversals, syllogistic inferences, and other logical deductions.
  • We propose to bridge this gap via dataset augmentation—prompting an LM to reason about the data in-context, and then adding these reasoning traces to the training set.
@@
Reversal curse paper
We first use the reversal dataset proposed by berglund2024reversal that contains one-sentence descriptions of fictional celebrities. Dataset examples can have the celebrity name (e.g. ‘Daphne Barrington’) preceding the description (e.g. ‘the director of "A Journey Through Time."’) or vice-versa. The dataset includes two independent sets of celebrities – ‘A’ and ‘B’. Following berglund2024reversal, we finetune the model on celebrities from the set ‘A’ with names preceding the description, and from set ‘B’ where names and descriptions appear in both orders but in separate examples. Overall, the train set consists of 3600 examples. The test set evaluates whether the model can infer the celebrity names from the set ‘A’ given only their descriptions. To add distractors in test examples, we include three randomly selected incorrect celebrity names in the list of options to score.
@@
We then test our methods with two simple datasets [^2] These datasets are adapted from the nonsense NLI and Syllogism datasets from lampinen2024language. that contain independent examples of reversal and syllogistic reasoning.
Simple reversals: Each training example consists of a brief preamble (e.g. “Did you know that”) followed by a single statement comparing two entities, e.g. “femp are more dangerous than glon.” There are a hundred such comparisons provided (with the comparison words sampled from a set of 28, e.g. “brighter”, “heavier,”). Each comparison is repeated across 10 different training examples, each time paired with a randomly sampled preamble. The test set consists of forced choices between the correct reversal and a contradictory relation, e.g.: “glon are less/more dangerous than femp.” Note that the incorrect answer is a tempting alternative, as it uses the same words as the trained statement (just reordered), which means that a model relying on simple features like bag-of-words would tend to prefer the incorrect answer.
Simple syllogisms: There are 69 training examples. Each example consists of a preamble (which specifies the entities under discussion), followed by two statements that form a syllogism. For example: “The relationships between glon, troff and yomp are: All glon are yomp. All troff are glon.” The test examples evaluate whether the model makes a correct inference from that syllogistic form. They provide a preamble specifying two entities from the training example, and then score all possible statements about the relationship between the two entities involving the quantifiers and relations in the dataset. Following lampinen2024language, we omit the partial negative form (some X are not Y); thus, there are six such possible statements (the product of “all”, “some”, and “no” together with the two directions each of those relations could take). For example, the preamble for the test example corresponding to the syllogism above would be “The relationships between yomp and troff are:” and the correct answer would be “All troff are yomp.”
@@
Figure 1: A semantic structure with a hierarchy of properties and relations. (Reproduced with permission from rogers2008precis.)
The benchmark is built around a relational semantic hierarchy that allows deductive inferences and abstractions. This hierarchy is based on real-world categories and relations, and draws inspiration from prior work in cognitive science that has studied learning of rich semantic hierarchies (see Fig. 1 for an intuitive example that composes part of the real-world structure). We similarly create a hierarchical knowledge structure involving 110 categories of animals and objects, and their properties and relations. Unlike the original structure, it is not balanced in branching factors or depth, and properties are similarly asymmetrically distributed across entities (1-6 per category, plus inherited ones)—making its distributional statistics more naturalistic (see Appx. A.2 for details). This underlying structure is derived from the real world.
@@
The use of nonsense terms potentially poses tokenization challenges; however, we ameliorate these by generating short 4-5 letter nonsense words using plausible combinations of phonemes for English (sampled via gao2023scope). Moreover, we show in the next sections that in practice the models can readily make inferences about these nonsensical entities in context, so use of unfamiliar words is not the primary bottleneck on model performance.
@@
Test sets: Within the semantic structure, our test sets accommodate reversals of relations (gruds [dogs] are a type of abmes [mammals] => abmes include gruds; cf. berglund2024reversal), syllogism-like deductive inferences (e.g. gruds [dogs] are abmes [mammals]; abmes are rony [warm-blooded] => gruds are rony), and longer deductions. Specifically, we focus on the following splits (in rough difficulty order):
@@
  • Syllogisms over trained facts; that is, statements that can be logically inferred by linking two (implicitly) quantified statements in the training corpus, using an entity that overlaps between them.
  • Category holdouts: only one fact about a category is present in training: what its parent category is. All possible inferences from that fact are tested. This overlaps with some aspects of the syllogism splits, except that: 1) the information about the target category is more strictly limited, thus limiting other cues that could aid generalization; and 2) inferring the fact may require chaining more than two statements together.
@@
We evaluate the models both with and without few-shot QA examples taken from the validation set in the prompt; we find relatively similar results across these conditions so we collapse across them in our main analyses.
@@ Methods
We perform evaluation using multiple-choice likelihood scoring. We do not provide the answer choices in context. We generally construct the datasets to provide challenging alternative answers (see above).
In-context evaluation: To perform full-dataset in-context evaluation, we concatenate the documents in a training dataset and provide them as context to the (instruction-tuned) model. We then prompt the model to use that context to answer the question. We randomly subsample the documents by a factor of 8x when evaluating models in-context on the largest datasets, as there is some redundancy in the datasets, and we find that models experience interference as the context length scales.
@@
Our fine-tuning experiments mainly involve tuning Gemini 1.5 Flash (team2024gemini) on our datasets. We generally fine-tune with batch size 8 or 16 and learning rate $3\cdot 10^{-4}$ for 200-1000 steps (depending on dataset size and loss).
@@
Our key approach to dataset augmentation is to take advantage of the generalization of models during in-context reasoning in order to improve the coverage of the fine-tuning dataset. We approach this via several methods, but all chiefly aim at the goal of spending training time compute for in-context inference to create more finetuning data, in order to improve generalization out-of-context (i.e., without requiring additional information in context) at test time.
@@
Some datasets, such as the fictional celebrities dataset by berglund2024reversal and our semantic structure dataset, contain documents comprising multiple sentences that are logically or semantically linked. We find that splitting these documents at sentence-level, into multiple finetuning examples, leads to a much improved finetuning performance – even after accounting for the total dataset size and gradient steps. We explore two ways of splitting a document: 1) Independent Splitting: wherein all $n$ sentences of a document are split independently into $n$ separate training examples, 2) Cumulative Splitting: wherein an $n$-sentence document is split into $n$ examples, with the $i$th example contains all sentences up to the $i$th sentence. We analyze the effect of sentence splitting on the model generalization in appendix B.1. In the following sections, we assume independent sentence splitting unless stated otherwise.
@@
In Fig. 2, we first explore generalization in the context of the reversal curse phenomenon and dataset released by berglund2024reversal. We replicate the finding that finetuning on the forward directions does not produce generalization to the reversals. The authors of that work noted briefly that in-context the models could perform better at this task. We study this more systematically by presenting the entire dataset in context, and find that the model performs nearly at ceiling on the reversals—thus providing a strong demonstration of the benefits of in-context learning over finetuning. Finetuning with data augmented with in-context inferences yields similarly high test performance. Simple finetuning, on the other hand, has near zero accuracy as the finetuned model always prefers those (incorrect) celebrity name completions that were seen as targets during training, regardless of the context; both ICL and Augmented Finetuning significantly outpeform base finetuning (both $t>149$, $p<10^{-16}$). Finally, a pretrained model performs near chance on the test set, indicating a lack of contamination.
@@
Simple nonsense reversals: We then correspondingly test the models on the simple nonsense version of our reversal dataset (Fig 3, left). We find a weaker, but still noticeable, benefit of ICL over finetuning in this setting, and again stronger benefits of augmented finetuning (respectively $ts=3.36,7.25$, $ps<10^{-3},10^{-11}$). The difference in benefits compared to the experiments above may be due to differences in the plausibility of the relationships in question, e.g. the possibility that nonsense words interfere to some degree with the model’s reasoning over longer contexts (see Appx. B.2).
Simple syllogisms: We next test the simple syllogisms dataset (Fig 3, right). Again, the pretrained model performs at chance, indicating a lack of contamination. Finetuning on the dataset does yield some non-trivial above-chance generalization; perhaps because for certain flavors of syllogism the logical inferences are compatible with simpler linguistic patterns — e.g., tuning on sequences like “All X include Y.” “All Y include Z.” might make the model more likely to predict “All X include Z” from pure associations. However, following simpler associative patterns is not valid for most syllogistic forms, e.g. if the universal quantifier is replaced by the existential in the examples above. Perhaps because of this, in-context learning yields much stronger performance; using in-context learning to augment the training dataset further improves performance overall (respectively $ts=2.83,6.30$, $ps<0.01,10^{-8}$).
@@
We then test the semantic structure benchmark (Fig. 4), which integrates multiple kinds of generalization in a richer learning setting. In this setting, we evaluate performance on: a) rephrased training statements (that preserve the direction of the relation) denoted as ‘train’ in the figure, b) reversals, c) syllogistic inferences, and d) propositions about held out categories. Across these settings, we find an overall benefit of ICL over finetuning, though the magnitude of this benefit varies depending on the split. We find some improvements in generalization to even rephrased information from training, and more dramatic improvements on reversals and syllogisms. However, the category-holdout split remains difficult, and improvements from ICL are minimal. Furthermore, we continue to find that augmenting the finetuning data with in-context inferences improves performance from finetuning, in many cases outperforming even ICL. Most improvements are statistically significant ($ts>2.79$, $p<0.01$) except for ICL on Syllogisms and both on Category show only marginal improvements ($1.95>ts>1.93$, $p=0.05$). (N.B., in this setting we do not use sentence splitting for the finetuning baseline, as we find it impairs performance and we want to compare to the strongest baselines; results with sentence-splitting for the baseline, along with other ablations, can be found in Appx. B.5.)
Our results in this section also highlight an important nuance to prior results on the reversal curse (berglund2024reversal). When the information being tested forms part of a broader coherent structure of knowledge (such as our semantic structure), finetuning alone does exhibit some above-chance generalization to reversals. This generalization may be due to the fact that other information in the training set can support the reversed conclusion; e.g. if the reversal were from “birds include eagles” to “eagles are a type of bird,” but the training set also includes statements such as “eagles have wings,” that information may enable an alternative route to inferring the reversal (associatively if not logically). Nevertheless, in-context learning and augmented finetuning continue to perform substantially better than finetuning alone in most cases.
@@
Process knowledge: We also performed some preliminary experiments exploring learning and generalization of process-type (rather than semantic) knowledge, i.e. the ability to apply novel procedures to inputs. Specifically, we focused on executing a simple pseudo-mathematical procedure that we call “derivatoids”, which transforms math-like expressions in a manner similar to computing a derivative. We created a train-test compositional generalization split in which certain combinations of rules are seen in training (again, either in context or via finetuning), but other combinations are held out for testing. The overall pattern of results is generally consistent with those above. ICL outperforms finetuning, especially in the low-shot regime, and augmentations improve finetuning performance. However, process knowledge requires distinct evaluation methods, and the effects correspondingly appear to be driven by distinct factors compared to the semantic knowledge experiments above; for brevity, we present the experiments and results in full detail in Appendix B.6.
@@
In-context learning: A variety of works have explored in-context learning (brown2020language; dong2022survey; lampinen2024broader). Several works have studied patterns in the learning and generalization of ICL, either empirically (garg2022can; chan2022data), mechanistically (olsson2022context), or theoretically (xie2023explanation; zhang2024trained; elmoznino2024context). Several recent works have found that even scaling to hundreds or thousands of in-context examples can improve LM performance on challenging datasets (agarwal2024many; anil2024many; bertsch2024context). Our work similarly relies on the ability of models to learn from many documents in context in order to generate the inferences for augmentation. Our work differs from prior works in focusing more on generalization of broader types of information in context, beyond few-shot tasks (cf. lampinen2024broader).
@@
Out-of-context learning: Several other papers (e.g. berglund2023taken; meinke2023tell) have considered how models can generalize “out-of-context”—that is, how they can use information that is not directly included in the prompt in a flexible way at test time. Our results may connect to these, e.g. in the patterns of partial generalization seen even with finetuning on the semantic structure dataset. However, we generally do not find reliable use of out-of-context information as in-context—i.e., in-context learning tends to outperform finetuning.
@@
Factual learning & generalization: A variety of works have studied how language models learn and generalize factual information, both in pretraining (e.g. allen2024physics3_1; allen2025physics3_2; zucchet2025language) and during finetuning (e.g. berglund2024reversal; gekhman2024does). Many of these works have focused on failures to generalize like the “Reversal Curse” (berglund2024reversal), while others looked into increases in hallucination as LLMs are pretrained (zucchet2025language) or finetuned on new knowledge (gekhman2024does). However, these works have generally not performed data-matched experiments comparing the generalization of in-context learning and finetuning.
@@
Data augmentation: A wide variety of works have explored how LLMs can be used to augment data for improved performance from small or narrow datasets, e.g. to improve zero-shot task performance (maini2024rephrasing) or to improve cross-lingual generalization (e.g. whitehouse2023llm). ding2024data reviews some of this broader literature. There have also been targeted attempts to fix specific issues like the reversal curse with hardcoded augmentations (golovneva2024reverse). A closely related work by akyurek2024deductive proposes “deductive closure training” to improve coverage via prompting language models to generate deductive inferences from training documents. padmanabhan2024propagating similarly propose generating continuations and then distilling them into the model. Several concurrent works chen2024reverse; ruan2025reasoning suggest that having LMs generate additional reasoning directions, and using these to augment their training data, can improve performance on reasoning tasks. kujanpaa2024knowledge propose prompt distillation, which distills from larger prompts containing privileged information, and find that this improves generalization from knowledge in finetuning. yang2024synthetic use language models to extract entities in the training documents, and generate synthetic data reasoning about links between these entities; like our results, they find this method of “rearranging” knowledge in synthetic data helps to improve downstream performance. Finally, a concurrent work (park2025textit) finds similar benefits to ours of in-context learning over finetuning, and likewise proposes augmenting data to overcome it. Our experiments show that in controlled settings, without the possibility of dataset contamination, similar approaches to augmenting small finetuning datasets can yield improved generalization—and relate the performance to that achieved through more basic finetuning and in-context learning.
@@
Synthetic data: An equally broad literature has explored other applications of synthetic data to improving LM performance; see liu2024best for a recent survey. Earlier works considered hand-designed synthetic data targeted to improve generalization using domain knowledge of areas like linguistics or mathematics (e.g. pratapa2018language; wu2021lime). More recent approaches have focused on generating the data directly from language models, either filtering with ground-truth scores (zelikman2022star) or simply through self-consistency (wang2023self; huang2023large). While a recent prominent article argued that training on synthetic data leads models to collapse (shumailov2024ai), other works have noted that in settings where synthetic data is mixed with the original data performance continues to improve (gerstgrasser2024model). We correspondingly find that in our setting, rather than being harmful, synthesizing augmented data with the models consistently improves performance (even on rephrased information from the training split in the semantic structure dataset). These results contribute to the ongoing discussion of how incorporating synthetic data affects model performance.